Tom Johnson and Melissa Acino started driving at the same time in different cars from Oklahoma City. They both traveld south on Route 35. Whem melissa covered a distance of 150 miles, tom had only covered a distance of 122 miles. If melissa avveraged 14 miles per hour faster than tom, find the average speed of each car.
Answer by mananth(16949)

Tom Johnson and Melissa Acino started driving at the same time in different cars from Oklahoma City. They both traveld south on Route 35. Whem melissa covered a distance of 150 miles, tom had only covered a distance of 122 miles. If melissa avveraged 14 miles per hour faster than tom, find the average speed of each car.
let tom.s rate be x mph
melissa's rate will be x+14 mph
Time = Distance / rate
150 / x+14 = 122 / x ( same time they cover different distances)
122(x+14) = 150 x
122x + 1708 =150x
28x= 1708
x= 1708 / 28
61 mph that is tom"s rate

Ananth
61+14 = 75 mph is Melissa's rate